OpenStudy (anonymous):

f(x,y)=|x|+|y| for all (x,y) belonging to R^2 Show that for the above function lim (x,y)->(0,0) for f(x,y) , exists using epsilon-delta method.

OpenStudy (anonymous):

Let \[ \epsilon >0\] and let \[ \delta =\frac {\epsilon}{2} \] If \[ |x| \le \delta , \quad |y| \le \delta\] then \[| f(x,y)- f(0,0)| =|x|+|y| \le 2 \delta =\epsilon \]